dy/dx +[(x 2 -1)(y 2 -1)] 1/2 /xy=0

dy/dx +[(x2-1)(y2-1)]1/2/xy=0

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear upendra

dy/dx +[(x2-1)(y2-1)]1/2/xy=0

or  ydy/(y2-1)1/2      =    - dx (x2-1)1/2/x

or ydy/(y2-1)1/2      =    - dx (x2-1)/x(x2-1)1/2

or ydy/(y2-1)1/2      =   x dx /x2(x2-1)1/2     -xdx /(x2-1)1/2    

 

now intigrate

(y2-1)1/2    = tan-1(x2-1)1/2   - (x2-1)1/2    


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin